OpenStudy (anonymous):

Solve the inequality for y. y - 3x > -5 HELP HELP HELP HELP

OpenStudy (kropot72):

If you add 3x to both sides of the inequality we get: y - 3x + 3x = 3x - 5 Can you simplify the left hand side?

OpenStudy (anonymous):

y>3x-5

OpenStudy (anonymous):

you dont flip the sign or anything

OpenStudy (kropot72):

You are correct. Sorry I left out the greater than symbol in my equation.

OpenStudy (anonymous):

so do you flip the sign to >< or keep it the same

OpenStudy (kropot72):

No, you don't reverse the symbol in this case, the reason being we are not multiplying both sides by -1.

OpenStudy (kropot72):

Does that make sense?

OpenStudy (anonymous):

What is the value of y when x = -3? 4x + 2y = 8 is it 2

OpenStudy (kropot72):

Not really. 4x + 2y = 8 Plugging in x = -3 we get: \[(4\times-3)+2y=8\] Therefore -12 + 2y = 8 Now can you add 12 to both sides? What do you get?